Rinehart Corporation purchased from its stockholders 5,000 shares of its own previously issued stock for $255,000. It later resold 2,000 shares for $54 per share, then 2,000 more shares for $49 per share, and finally 1,000 shares for $43 per share.
Prepare journal entries for the purchase of the treasury stock and the three sales of treasury stock. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Treasury Stock 255000 Cash 255000 (To record purchase from stockholders.) Cash 108000 Treasury Stock Paid-in Capital from Treas To record sales of shares at $54 per share.) Cash Paid -in Capital rom Treast 98000 Paid-in Capital from Treasury Stock (To record sales of shares at $49 per share.) Treas Cash Paid-in Capital from Treas
Answer #1
Concepts and reason

Journal Entry: This is the primary recording of any transaction which is entered in the financial accounting. First of all, every transaction is put into a journal entry and from there it is posted to different accounts. Journal entry forms the basis of accounting. If a journal entry is wrong then financial books cannot provide a correct picture of the financials of the entity.

Fundamentals

Treasury stock: This is that part of the common stock of the company that is re-purchased by the issuing company. In other words, these are shares that were initially issued but later on, the issuing company bought them in order to reduce its outstanding stock.

Additional paid-in the capital: It is that value that is paid by the holder of the common stock that is over and above the face value of the share. It is shown in the balance sheet in the shareholder’s equity section.

Retained earnings: These are the amount which is left after distribution of dividends and proposed dividends to shareholders of the company from the opening balance of retained earnings and net income of the year.

Rules of debit and credit have been followed for journalizing the various transactions and these are as mentioned below:

Debit the Receiver and Credit the Giver: It is used for personal accounts. It indicates when the organization receives something from anyone then, what is received would be debited and the giver would be credited.

Debit what comes in and Credit what goes out: It is used for real accounts. It indicates when the organization purchased or receives any asset then it would be debited and on the other side, when the asset is going out of the organization then, it would be credited.

Debit all expenses and losses, credit all incomes and gains: It is used in case of a nominal account, and according to this rule all the expenses and losses incurred by the organization would be debited and on the other side, all the incomes and gain of the organization would be credited.

Pass the Journal entry using the MS-Excel to record the purchase of Treasury stock:

Record the Journal entry using the MS-Excel to record the sale of treasury stock:

Working notes:

Compute the cash received from the sale of treasury stock using the equation as shown below:

Totalcashreceived=Numberofshares×Pricepershares=2,000×$54=$108,000\begin{array}{c}\\{\rm{Total cash received}} = {\rm{Number of shares}} \times {\rm{Price per shares}}\\\\ = 2,000 \times \$ 54\\\\ = \$ 108,000\\\end{array}

Hence, the total cash received is $108,000.

Compute the value of treasury stock sold using the equation as shown below:

TreasuryStock=PurchasepriceNumberofsharespurchase×Numberofsharessold=$255,0005,000×2,000=$102,000\begin{array}{c}\\{\rm{Treasury Stock}} = \frac{{{\rm{Purchase price}}}}{{{\rm{Number of shares purchase}}}} \times {\rm{Number of shares sold}}\\\\ = \frac{{\$ 255,000}}{{5,000}} \times 2,000\\\\ = \$ 102,000\\\end{array}

Hence, the value of treasury stock sold is $102,000.

Compute the paid in capital from treasury stock using the equation as shown below:

Paidincapital=(SellingpriceoftreasurystockPurchasepriseoftreasurystock)=$108,000$102,000=$6,000\begin{array}{c}\\{\rm{Paid in capital}} = \left( \begin{array}{c}\\{\rm{Selling price of treasury stock}} - \\\\{\rm{Purchase prise of treasury stock}}\\\end{array} \right)\\\\ = \$ 108,000 - \$ 102,000\\\\ = \$ 6,000\\\end{array}

Hence, paid in capital is $6,000.

Record the Journal entry using the MS-Excel to record the sale of treasury stock:

Working notes:

Compute the cash received from the sale of treasury stock using the equation as shown below:

Totalcashreceived=Numberofshares×Pricepershares=2,000×$49=$98,000\begin{array}{c}\\{\rm{Total cash received}} = {\rm{Number of shares}} \times {\rm{Price per shares}}\\\\ = 2,000 \times \$ 49\\\\ = \$ 98,000\\\end{array}

Hence, total cash received is $98,000.

Compute the value of treasury stock sold using the equation as shown below:

TreasuryStock=PurchasepriceNumberofsharespurchase×Numberofsharessold=$255,0005,000×2,000=$102,000\begin{array}{c}\\{\rm{Treasury Stock}} = \frac{{{\rm{Purchase price}}}}{{{\rm{Number of shares purchase}}}} \times {\rm{Number of shares sold}}\\\\ = \frac{{\$ 255,000}}{{5,000}} \times 2,000\\\\ = \$ 102,000\\\end{array}

Hence, the value of treasury stock sold is $102,000.

Record the Journal entry using the MS-Excel to record the sale of treasury stock:

Working notes:

Compute the cash received from the sale of treasury stock using the equation as shown below:

Totalcashreceived=Numberofshares×Pricepershares=1,000×$43=$43,000\begin{array}{c}\\{\rm{Total cash received}} = {\rm{Number of shares}} \times {\rm{Price per shares}}\\\\ = 1,000 \times \$ 43\\\\ = \$ 43,000\\\end{array}

Hence, the total cash received is $43,000.

Compute the value of treasury stock sold using the equation as shown below:

TreasuryStock=PurchasepriceNumberofsharespurchase×Numberofsharessold=$255,0005,000×1,000=$51,000\begin{array}{c}\\{\rm{Treasury Stock}} = \frac{{{\rm{Purchase price}}}}{{{\rm{Number of shares purchase}}}} \times {\rm{Number of shares sold}}\\\\ = \frac{{\$ 255,000}}{{5,000}} \times 1,000\\\\ = \$ 51,000\\\end{array}

Hence, the value of treasury stock sold is $51,000.

Compute the amount charged from retained earnings using the equation as shown below:

Amountchargedfromretainedearnings=(PurchasepriceSalespricePaidincapitalfromtreasurystock)=$51,000$43,000$2,000=$6,000\begin{array}{c}\\{\rm{Amount charged from retained earnings}} = \left( \begin{array}{c}\\{\rm{Purchase price}} - {\rm{Sales price}}\\\\ - {\rm{Paid in capital from treasury stock}}\\\end{array} \right)\\\\ = \$ 51,000 - \$ 43,000 - \$ 2,000\\\\ = \$ 6,000\\\end{array}

Hence, the amount charged from retained earnings is $6,000.

Ans: